OpenStudy (anonymous):

ndicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5

OpenStudy (anonymous):

y = m x + b Apply the values from the problem statement. -2 = (2/5)5 + b b = - 4 Slope intercept of the line is: y = (2/5)x - 4 Standard form: Move the x term to the RHS. - (2/5)x + y = -4 Multiply each side of the above by -1 so that the coefficient of x is positive.\[\frac{2 x}{5}-y=4 \]

OpenStudy (anonymous):

Sorry. A mistake in a sentence above. "Move the x term to the RHS." should have been written: Move the x term to the LHS.